# Berry phases and connection matrices defined on homogeneous spaces attached to Jacobi groups

### v2

@@ 4.1. Definitions
We endow the $n(n\!+\!3)\!+\!1$-dimensional manifold $\tilde{{\mathcal{X}}}^{J}_{n}$ with an ACOS structure $(\tilde{{\mathcal{X}}}^{J}_{n},\theta,\omega)$. For $\theta$ we take the formula (4.1), and consider $\Omega$ written in Darboux coordinates
@@
[…]
@@ 4.3. Covariant derivative of one-forms on ${\mathcal{X}}^{J}_{1}$ and $\tilde{{\mathcal{X}}}^{J}_{1}$
(4.24) $-D(\operatorname{d}w)=\left(\begin{array}[]{cc}\operatorname{d}z\operatorname{d}w\end{array}\right)\left(\begin{array}[]{cc}\lambda&\lambda\bar{\eta}\\ \lambda\bar{\eta}&\lambda\bar{\eta}^{2}+2\frac{\bar{w}}{P}\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}z\\ \operatorname{d}w\end{array}\right).$
@@ 5.1.1. Connections on real manifolds
(5.1a) $1$ $~{}~{}D(s_{1}+s_{2})=Ds_{1}+Ds_{2},\quad\forall s_{1},s_{2}\in\Gamma(E),$
(5.1b) $2$ $~{}~{}D(\alpha s)=\operatorname{d}\alpha\otimes s+\alpha Ds,\quad\alpha\in C^{\infty}(M).$
@@
where <,><,>< , > is the pairing between the tangent space $T(M)$ and the cotangent space $T^{*}(M)$.
@@
(5.20a) $1$ $\nabla_{fX+gY}=f\nabla_{X}+g\nabla_{Y},~{}X,Y\in{\mathfrak{D}}^{1},~{}f,g\in C^{\infty}(M),$
(5.20b) $2$ $\nabla_{X}(fY)=f\nabla_{X}(Y)+(Xf)Y,\quad Xf=<X,\operatorname{d}f>.$
@@
where $\omega^{\prime j}_{i}=\Gamma^{\prime j}_{ik}\operatorname{d}\omega^{k}$. Fomula (5.23c) appears also as [60, (2) page 27]. $DX$ is a (1,1)-type tensor field on $M$, called the absolute differential of $X$, [49, page 114].
Let $T^{r}_{s}$ be the tensor product of the tangent and cotangent bundles. If $t$ is an $(r,s)$-type tensor field, then the image of $t$ under the induced connection $D$ is an $(r,s+1)$-type tensor field $Dt$.
@@ 5.1.2. Connections on complex manifolds
Several definitions introduced in §5.1.1 for real connections are adapted to the complex case [48]. Essentially, the transpose of $A^{t},~{}A\in M(q,\mathbb{R})$ is replaced by the conjugate transpose (or hermitian transpose) $A^{H}$, $A\in M(q,\mathbb{C})$, also denoted $A^{*}$, $A^{\dagger}$, $A^{+}$.
@@
Using the decomposition $T^{*}_{M}=(T_{M}^{*})^{\prime}+(T_{M}^{*})"$, $D=D^{\prime}+D"$, where $D^{\prime}:{\mathfrak{A}}^{0}(E)\rightarrow{\mathfrak{A}}^{1,0}$ and $D":{\mathfrak{A}}^{0}(E)\rightarrow{\mathfrak{A}}^{0,1}$. The connection $D$ is called compatible with the complex structure if $D"=\bar{\partial}$ [59, page 73].
@@
(5.35a) $1$ $~{}~{}H(\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2},\eta)=\lambda_{1}H(\xi_{1},\eta)+\lambda_{2}H(\xi_{2},\eta),\quad\lambda_{1},\lambda_{2}\in\mathbb{C},~{}~{}\xi_{1},\xi_{2},\eta\in V;$
(5.35b) $2$ $~{}~{}\overline{H(\xi,\eta)}=H(\eta,\xi),\quad\xi,\eta\in V.$
@@
and we find for the tautological line bundle $[-1]$ on ${\mathbb{CP}}^{n}$
@@
$\Theta_{[1]}=\frac{(1+||w||^{2})\sum\operatorname{d}w_{k}\wedge\operatorname{d}\bar{w}_{k}-\sum\bar{w}_{k}\operatorname{d}w_{k}\wedge\sum w_{k}\operatorname{d}\bar{w}_{k}}{(1+||w||^{2})^{2}}.$
@@
In the convention $\alpha,\beta,\gamma,\dots$ run from 1 to $n$, while A,B,C,$\dots$ run through $1,\dots,n,$ $\bar{1},\dots,\bar{n}$, [64, p 155], for an almost complex connection without torsion we have the relations
@@
To a Kähler manifold $(M,\omega_{M})$ it is attached the triple $({L},h,\nabla_{L})$ [81, 37], where ${L}$ is a holomorphic (prequantum) line bundle on $M$, $h$ is a hermitian metric on ${L}$ (taken conjugate linear in the first argument), and $\nabla_{L}$ is a connection compatible with the metric and the Kähler  metric,
@@
Then $\omega_{M}$ is integral, i.e. $\omega_{M}\in H^{2}(M,\mathbb{Z})$ and the first Chen class $c_{1}[{L}]=[\omega_{M}]$ [59, page 141]. $M$ is called a Hodge manifold for compact Kähler  manifolds. Then [][][ ]
$[]:\operatorname{Div}(M)\rightarrow(M,{\mathcal{O}}^{*})$
@@ 5.2. Examples
Glauber’s coherent states $e_{z}=e^{z{\mbox{\boldmath{$a$}}}^{\dagger}}e_{0}$ have the scalar product
@@
[…]
@@
$e_{z}=e^{z{\mbox{\boldmath{$K$}}}_{+}}e_{k,k},\quad|z|<1,\quad k=1,\frac{3}{2},2,\frac{5}{2},\dots,$
@@
Let us denote with $J$ the extremal weight of the representation (5.55) ((5.57)) i.e. $J=-2j$ $(J=2k)$ of $\text{SU}(2)$ (respectively, $\text{SU}(1,1)$). Then we write (5.58) as
@@
[…]
@@
We recall also the definition of the tautological line bundle $[-1]$
[…]
@@
$[1]$ is the hyperplane bundle, the dual of the tautological bundle $[-1]$.

### v3

@@ 4.1. Definitions
We endow the n(n+3)+1n(n\!+\!3)\!+\!1-dimensional manifold $\tilde{{\mathcal{X}}}^{J}_{n}$ with an ACOS structure $(\tilde{{\mathcal{X}}}^{J}_{n},\theta,\omega)$. For $\theta$ we take the formula (4.1), and consider $\Omega$ written in Darboux coordinates
@@
Finally, $\omega_{{\mathcal{X}}^{J}_{n}}(x,y,q,p)$ corresponds to $\Omega$ $(Q^{I},P^{I}),I=1,\dots,N_{1}$ in (4.8) respectively $i=1,\dots,n$ in (4.9).
@@ 4.3. Covariant derivative of one-forms on ${\mathcal{X}}^{J}_{1}$ and $\tilde{{\mathcal{X}}}^{J}_{1}$
(4.24) $D(\operatorname{d}w)=\left(\begin{array}[]{cc}\operatorname{d}z\operatorname{d}w\end{array}\right)\left(\begin{array}[]{cc}\lambda&\lambda\bar{\eta}\\ \lambda\bar{\eta}&\lambda\bar{\eta}^{2}+2\frac{\bar{w}}{P}\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}z\\ \operatorname{d}w\end{array}\right).$
@@ 5.1.1. Connections on real manifolds
(5.1a) $1.$ $~~D(s_{1}+s_{2})=Ds_{1}+Ds_{2},\quad\forall s_{1},s_{2}\in\Gamma(E),$
(5.1b) $2.$ $~~D(\alpha s)=\operatorname{d}\alpha\otimes s+\alpha Ds,\quad\alpha\in C^{\infty}(M).$
@@
where <,><,> is the pairing between the tangent space $T(M)$ and the cotangent space $T^{*}(M)$.
@@
(5.20a) $1.$ $\nabla_{fX+gY}=f\nabla_{X}+g\nabla_{Y},~X,Y\in{\mathfrak{D}}^{1},~f,g\in C^{\infty}(M),$
(5.20b) $2.$ $\nabla_{X}(fY)=f\nabla_{X}(Y)+(Xf)Y,\quad Xf=<X,\operatorname{d}f>.$
@@
where $\omega^{\prime j}_{i}=\Gamma^{\prime j}_{ik}\operatorname{d}\omega^{k}$. Fomula (5.23c) appears also as [60, (2) page 27]. DXDX is a (1,1)-type tensor field on $M$, called the absolute differential of $X$, [49, page 114].
Let $T^{r}_{s}$ be the tensor product of the tangent and cotangent bundles. If $t$ is an $(r,s)$-type tensor field, then the image of $t$ under the induced connection $D$ is an $(r,s+1)$-type tensor field DtDt.
@@ 5.1.2. Connections on complex manifolds
Several definitions introduced in §5.1.1 for real connections are adapted to the complex case [48]. Essentially, the transpose of $A^{t},~A\in M(q,\mathbb{R})$ is replaced by the conjugate transpose (or hermitian transpose) $A^{H}$, $A\in M(q,\mathbb{C})$, also denoted $^{*}, AA^{\dagger}$, $A^{+}$.
@@
[…]
@@
(5.35a) $1.$ $~~H(\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2},\eta)=\lambda_{1}H(\xi_{1},\eta)+\lambda_{2}H(\xi_{2},\eta),\quad\lambda_{1},\lambda_{2}\in\mathbb{C},~~\xi_{1},\xi_{2},\eta\in V;$
(5.35b) $2.$ $~~\overline{H(\xi,\eta)}=H(\eta,\xi),\quad\xi,\eta\in V.$
@@
and we find for the tautological line bundle [1][-1] on ${\mathbb{CP}}^{n}$
@@
[…]
@@
[…]
@@
To a Kähler manifold $(M,\omega_{M})$ it is attached the triple $({L},h,\nabla_{L})$ [81, 37], where L{L} is a holomorphic (prequantum) line bundle on $M$, $h$ is a hermitian metric on L{L} (taken conjugate linear in the first argument), and $\nabla_{L}$ is a connection compatible with the metric and the Kähler  metric,
@@
Then $\omega_{M}$ is integral, i.e. $\omega_{M}\in H^{2}(M,\mathbb{Z})$ and the first Chen class $c_{1}[{L}]=[\omega_{M}]$ [59, page 141]. $M$ is called a Hodge manifold for compact Kähler  manifolds. Then $[]$
$\operatorname{Div}(M)\rightarrow(M,{\mathcal{O}}^{*})$
@@ 5.2. Examples
Glauber’s coherent states $e_{z}=e^{z{\mbox{{$a$}}}^{\dagger}}e_{0}$ have the scalar product
@@
(5.55) ${\mbox{{$J$}}}_{+}e_{j,-j}\not=0,\quad{\mbox{{$J$}}}_{-}e_{j,-j}=0,\quad{\mbox{{$J$}}}_{0}e_{j,-j}=-je_{j,-j},\quad j=\frac{m}{2},~m\in{\mathbb{N}}.$
@@
$e_{z}=e^{z{\mbox{{$K$}}}_{+}}e_{k,k},\quad|z|<1,\quad k=1,\frac{3}{2},2,\frac{5}{2},\dots,$
@@
Let us denote with $J$ the extremal weight of the representation (5.55) ((5.57)) i.e. J=2jJ=-2j $(J=2k)$ of $\text{SU}(2)$ (respectively, $\text{SU}(1,1)$). Then we write (5.58) as
@@
(5.60) $(e_{Z^{\prime}},e_{Z})=\det(\mathbb{1}_{n}+\epsilon ZZ^{\prime+})^{\epsilon},\text{~~where~~}\epsilon=~-~(+)\text{~~for~~}X_{n}~(X_{c}),$
@@
We recall also the definition of the tautological line bundle [1][-1]
(5.72) $1]=\{(z,v)\in{\mathbb{P}}(\mathbb{C}^{n+1})\times\mathbb{C}^{n+1}|~v\in[z],~[z]\text{~is the line bundle defined by z}\},$
@@
$[1]$ is the hyperplane bundle, the dual of the tautological bundle [1][-1].